OpenStudy (anonymous):

Solve the system by graphing.If the system is inconsistent and has no solution,state this .If the system is dependent,write the form of the solution for any real number x. 4x+2y=-8 y=3/4+7

OpenStudy (anonymous):

bring this to standard form, to graph it: 4x+2y=-8

OpenStudy (anonymous):

Ok let walk it together ...My problem is on how to graph

OpenStudy (anonymous):

Here is my solution now I need you to finish it for me 4x+2y=-8 y=(3/4)x+7 Now we rewrite as 4x+2y=-8 (3/4)x-y=-7 Divide the first equation by 2 and have 2x +y=-4 (3/4)x-y=-7 Add and get (2 3/4)x=-11 or (11/4)x=-11 or x=-4 and -16+2y=-8 gives y=4 (3/4)(-4)-4=-7

OpenStudy (anonymous):

it is easier to bring the equations to standard form if you need to graph them

OpenStudy (anonymous):

So can you help out and also graph then I will give medal

OpenStudy (anonymous):

4x+2y=-8 try to bring this to standard form y=

OpenStudy (anonymous):

y=-4

OpenStudy (anonymous):

4x + 2y = -8 2y = -8 - 4x y = -4 -2x

OpenStudy (anonymous):

ok? :)

OpenStudy (anonymous):

is that the final solution

OpenStudy (anonymous):

we now have two equations y = -4 -2x y=3/4+7

OpenStudy (anonymous):

we can graph them now

OpenStudy (anonymous):

Can you graph it

OpenStudy (anonymous):

yes wait a sec

OpenStudy (anonymous):

y=3/4x+7 <- 3/4 is x , ri ght?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you know y=3/4x+7 what each part of the standard form say about the line?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

ok, the number with no x is the shift upwards from the origin: y=x+3

OpenStudy (anonymous):

its called y-intercept "c" :)

OpenStudy (anonymous):

so is this solution inconsistent and has no solution or dependent

OpenStudy (anonymous):

we know if we draw both lines the x term is the slope or height of the line:

OpenStudy (anonymous):

ok?